OpenStudy (anonymous):

square root 24+ square root 54

OpenStudy (campbell_st):

\[\sqrt{24} = 2\sqrt{6}\] \[\sqrt{54} = 3\sqrt{6}\] the answer is \[5\sqrt{6}\]